### v1

# Investigation of unsupervised and supervised hyperspectral anomaly detection

Abstract
Hyperspectral sensing is a valuable tool for detecting anomalies and distinguishing between materials in a scene. Hyperspectral anomaly detection (HS-AD) helps characterize the captured scenes and separates them into anomaly and background classes. It is vital in agriculture, environment, and military applications such as RSTA (reconnaissance, surveillance, and target acquisition) missions. We previously designed an equal voting ensemble of hyperspectral unmixing and three unsupervised HS-AD algorithms. We later utilized a supervised classifier to determine the weights of a voting ensemble, creating a hybrid of heterogeneous unsupervised HS-AD algorithms with a supervised classifier in a model stacking, which improved detection accuracy. However, supervised classification methods usually fail to detect novel or unknown patterns that substantially deviate from those seen previously. In this work, we evaluate our technique and other supervised and unsupervised methods using general hyperspectral data to provide new insights.
keywords:
hyperspectral imaging, near-infrared NIR, remote sensing, unmanned aerial vehicles UAV, anomaly detection, machine learning, stacking ensemble learning
1 Introduction
Hyperspectral remote sensing is a technique that combines spectroscopy and imaging to collect spatial and spectral information and process information from across the electromagnetic spectrum. Hyperspectral imaging (HSI) aims to obtain the spectral information from a scene, and advancement of HSI technologies [1] combined with machine learning (ML) recognition techniques [2] promises improved scene characterization in numerous applications, such as the detection of anomalies in a large background, find objects, or identify materials. Hyperspectral anomaly detection (HS-AD) [3] helps characterize the captured scenes and separates them into anomaly and background classes. Supervised anomaly detection solves this problem using data labeled with ground truth, whereas unsupervised anomaly detection does not.
Universality or generalization suggests that if a method achieves promising results on known datasets, it will perform similarly well for new or unseen data and provide consistent predictions. Usually, unsupervised HS-AD methods promises generalization as it should work well enough on unknown data. Our previously introduced a new approach [4, 5] to the HS-AD problem used a supervised classifier as meta-model in a stacking [5] ensemble, which improved detection accuracy. However, we found that our previous method does not generalized well when trained on one dataset and tested on other unfamiliar dataset. Unlike supervised ML models, unsupervised HS-AD methods should work well enough on unknown data. Unfortunately, our evaluation found that unsupervised HS-AD methods are not consistently generalizing, and their performance varies from dataset to dataset.
In this treatment, we present our unsupervised stacking-ensemble anomaly detection approach to solve the generalization problem. We have utilized the Greedy search [5] approach to find suitable HS-AD methods for the base model and unsupervised Gaussian mixture model as the meta-model. We believe this approach holds significant potential for improving generalized anomaly detection accuracy. We have considered twenty-two scenes from four public and one private benchmarking datasets (thirteen scenes from the ABU (airport, beach, and urban) dataset, five from the Arizona dataset, two from the San Diego airport dataset, and one from each of the HYDICE urban and Salinas datasets). We used 2-fold cross-validation five times to show our methods’ statistical significance, where it achieved an average ROC-AUC score of 0.918 with a standard deviation of 0.131, which is at least seven % better than other individual methods. This article also investigates other supervised and unsupervised methods using these datasets. We compare the ROC-AUC scores to provide new insights and recommendations on the performance. Continuous improvement and innovation in this research area are crucial in further enhancing the general accuracy of hyperspectral anomaly detection and advancing the hyperspectral sensing field.
2 Background
Hyperspectral imagers are expensive, and only a few standard HS datasets are available [6]. As a result, most researchers are exploring unsupervised HS-AD algorithms [7] that do not require annotated data, such as those based on statistical methods [3].
2.1 Individual Hyperspectral Anomaly Detection Algorithms
The researchers and reviewers categorize hyperspectral anomaly detection (HS-AD) methods based on the design criteria to separate the background and anomalies and also based on the primary function model [8, 3, 7]. The Reed–Xiaoli (RX) [9, 10], MD–RX, and Windowed RX (WIN-RX) are some examples of Statistics-based HS-AD techniques. Subspace-RX (SSRX) [11], Complementary Subspace Detector (CSD) [12], Local Summation Anomaly Detection (LSAD) [13], Local Summation Unsupervised Nearest Regularized Subspace with an Outlier Removal Anomaly Detector (LSUNRSORAD) [14] are some examples of subspace-based techniques. Kernel–RX Algorithm (KRX) [15, 16], Gaussian Mixture RX (GM-RX) [17] are some examples of kernel-based techniques. Cluster-based anomaly detector (CBAD) [18], fuzzy c-means clustering-based anomaly detector (FCBAD) [19] are some examples of clustering distance-based techniques. Attribute and Edge-Preserving Filters (AED) [20] is an example of spatial-spectral filtering-based techniques. Isolation forest (iForest) [21] is an example of Machine learning-based techniques. Kernel iForest (KIFD) [22] combines both kernel-based and ML-based techniques. We found that each method’s performance depends upon specific background characteristic constraints and is effective only in some scenarios. Thus, we investigated to find a combination of methods to overcome this limitation.
2.2 Ensemble Hyperspectral Anomaly Detection Algorithms
Yang et al. [23] proposed an Ensemble and Random RX with Multiple Features (ERRX MF) anomaly detector using three features, Gabor, Extended Morphological Profile (EMP), and Extended Multiattribute Profile (EMAP), along with the original hyperspectral image. Wang et al. [24] proposed a sub-feature ensemble called SED. They randomly sub-sampled channels, and used six heterogeneous base models in the ensemble. Finally, a prior-based tensor approximation algorithm (PTA) [25] used these sub-features for anomaly detection. Fatemifar et al. [26] proposed a stacking ensemble for face spoofing anomaly detection that consisted of 63 base classifiers and a Gaussian Mixture Model (GMM) as the meta-classifier for the second stage. Younis et al. [4] proposed an equal-weight voting method called Hyperspectral Unmixing-Based Voting Ensemble Anomaly Detector (HUE-AD) that combines four detectors (Abundance, AED, KIFD, and LSUNRSORAD) to identify anomalies. Hossain et al. [5] proposed a stacking ensemble called Greedy Ensemble Hyperspectral Anomaly Detector (GE-AD) that uses a greedy search algorithm to find four suitable base detectors and utilizes a supervised machine learning model to combine those four base detectors to identify anomalies. Figure 1 shows the overall flowchart of the GE-AD algorithm. The greedy search algorithm uses 2-fold cross-validation to evaluate performance metrics and greedily searches for base methods to maximize the performance score, which reduces the search space compared to a grid search approach. This method produces a much higher F1-macro score and similar ROC-AUC score than the input methods.
Figure 1: Overall flowchart of GE-AD [5] to find the best AD methods for ensemble fusion.
3 Methodology
We optimized the HUE-AD and GE-AD in this treatment to improve their generalizability. We also investigated some unsupervised ML models as meta-models. Unlike the F1 score, we found that improving the ROC-AUC score in one dataset translates to a better ROC-AUC score in another. Thus, we tried to maximize the ROC-AUC score when we used greedy search. In this section, we discussed our methodology in detail.
3.1 Hyperspectral Unmixing-Based Voting Ensemble Anomaly Detector (HUE-AD)
Our empirical tests found that HUE-AD [4]’s three base anomaly detectors can operate adequately in a trio even when the particular spectra of interest or their abundances are unknown. Thus, under our new modification, we skip the abundance and determine a pixel as anomalous if at least two base methods vote a pixel as anomalous. Although not as good as the original HUE-AD, the performance was acceptable and still has a higher F1-macro score than the input methods. As it has a lower ROC-AUC score, we modified it further and used an average ensemble using the input results without binarizing to find the anomaly score. We are calling it the baseline in our investigation.
3.2 Greedy Ensemble Hyperspectral Anomaly Detector (GE-AD)
GE-AD [5] is a stacking ensemble that uses a supervised machine learning model as a meta-model that combines four detectors to identify anomalies. GE-AD utilizes a greedy search algorithm to find those four suitable detectors to maximize the scoring function. However, the trained model does not generalize and suffers when applied to an unfamiliar new scene. Thus, along with four base methods, we used thirty normalized random channels as input to the supervised meta-classifier. We expect the classifier to learn about the underlying dataset and generalize better. Our results show better ROC-AUC scores than before. We are calling this modified version the mGE-AD in our investigation.
3.3 Proposed Unsupervised Ensemble Approach (UGE-AD)
We have utilized the same greedy search [5] approach to find suitable HS-AD methods as the base model. We considered the Reed–Xiaoli (RX) [9, 10] detector, Attribute and Edge-Preserving Filters (AED) [20] , Cluster-based anomaly detector (CBAD) [18], Complementary Subspace Detector (CSD) [12], Fuzzy clustering-based anomaly detector (FCBAD) [19], Gaussian mixture RX (GM-RX) [17] anomaly detector, Kernel iForest (KIFD) [22] anomaly detector, Local Summation Unsupervised Nearest Regularized Subspace with an Outlier Removal Anomaly Detector (LSUNRSORAD) [14], Median RX (MD-RX) anomaly detector, Subspace-RX (SSRX) [11], Windowed RX (WIN-RX).
The RX [9, 10] detector has a simple statistical principle, low computational complexity, and relatively good performance. It characterizes background using the HSI’s mean and covariance, and calculates the Mahalanobis distance [27] between the background and pixel under test. As the mean is very susceptible to noise, MD-RX uses the median instead of the mean as statistical information to improve RX. Whereas, WIN-RX uses a sliding window to calculate local mean and covariance to determine anomalies. GM-RX [17] uses the Gaussian Mixture Model (GMM) to characterize the complex scene to improve the detection of the RX algorithm. Schaum et al. [11] in SSRX deleted several high-variance Principal Components (PC) dimensions as background clutters found in principal component analysis (PCA) to improve the RX performance. They later designed CSD [12] where the highest variance PCs are used to define the background subspace and the others (the complementary subspace) as the target subspace. Carlotto, Mark J. [18] proposed that the image background can be divided into clusters in the CBAD algorithm. FCBAD [19] improved the CBAD using fuzzy c-means clustering instead of k-means clustering. Kang et al. [20] assumed that anomalies tend to be smaller and possess unique reflectance signatures and that the pixels belonging to the same class would have a high correlation in the spatial domain, and based on this principle, they proposed AED. Tan et al. [14] proposed LSUNRSORAD to decrease the computational complexity of LSAD [13]. Li et al. [22] utilized the kernel space of PCA and the isolation forest (iForest) [21] to isolate anomalies in their proposed KIFD [22] algorithm.
We considered a similar idea to Wang et al. [24] and Fatemifar et al. [26] and investigated multiple unsupervised methods as meta-model. We used the Gaussian mixture model for the meta-model. We also investigated Forest; however, it did not perform well.
We have considered the proposed methodology of Yang et al. [23]. Instead of directly passing through the HSI data, we utilized PCA to find ten PCs as a passthrough to the meta-model to improve the ROC-AUC score.
3.4 Datasets
We assessed the AD algorithms’ effectiveness using twenty-two scenes from six public and one private benchmarking datasets. These scenes are from the Airport–Beach–Urban (ABU) dataset [20, 28], the San Diego dataset [29, 30], the Salinas dataset [31], the Hydice Urban dataset [32, 33], and the Arizona dataset [34].
The ABU airport and beach datasets have four scenes, whereas the ABU urban and the Arizona datasets have five. The San Diego airport dataset has two scenes. Both the HYDICE urban and the Salinas datasets have one scene each. Salinas dataset contains a farming field, whereas HYDICE urban has some urban development within vast vegetation. The ABU urban contains urban scenes. Arizona dataset contains a sandy desert scene with sparse vegetation and no urban development. The ABU beach contains scenes with water, which is challenging because of it. The ABU airport and the San Diego airport contain similar scenes from airports.
4 Results
We used the ROC-AUC (area under the receiver operating characteristic curve) [35, 36, 37] traditional classification evaluation metrics to evaluate the HS-AD algorithms.
For our first evaluation, we used only one dataset in the greedy search to find suitable base methods to maximize the ROC-AUC score. The results are shown in Table 1. The results show that the greedy search may fail to find the optimum base methods for small data, and the results are generalizing for other datasets except for the Arizona dataset. Table 1 shows that the performances are correlated to input base methods instead of the scene similarity. As we can see, GM-RX is required to get a suitable performance score for the San Diego airport dataset.
Table 1: Average ROC-AUC scores of our newly proposed Unsupervised ensemble UGE-AD over various datasets. The first column shows the dataset used in the greedy search to find suitable base methods, and the second column shows those base methods.
HYDICE Salinas San Diego ABU airport ABU beach ABU urban Arizona
HYDICE WIN-RX 0.997 0.951 0.962 0.943 0.963 0.964 0.802
Salinas LSUNRSORAD, RX 0.995 0.999 0.965 0.949 0.971 0.971 0.780
San Diego GM-RX 0.997 0.997 0.964 0.934 0.961 0.962 0.847
ABU airport AED, GM-RX, KIFD, LSUNRSORAD 0.995 1.000 0.971 0.962 0.985 0.966 0.712
ABU beach KIFD, LSUNRSORAD, RX, WIN-RX 0.996 0.997 0.965 0.960 0.973 0.976 0.766
ABU urban AED, KIFD, LSUNRSORAD, RX 0.996 1.000 0.970 0.962 0.985 0.971 0.706
Arizona CBAD, FCBAD, GM-RX, KIFD 0.994 0.999 0.972 0.948 0.967 0.966 0.779
For our second evaluation, from the twenty-two scenes from six public and one private benchmarking datasets, the GE-AD algorithm randomly selected datasets to ensure we have at least 50% image scenes are assigned to the greedy search. Over five runs with 2-fold cross-validation, greedy search tried to maximize the ROC-AUC score and selected AED, FCBAD, GM-RX, and KIFD as base methods for our unsupervised UGE-AD. In our unsupervised UGE-AD, we used these four base methods and ten PCs as input to the GMM meta-model. UGE-AD achieved an average of 0.918 ROC-AUC scores with a standard deviation (STD) of 0.131, which is lower than other input methods, as shown in Table 2 over seven datasets. The greedy search selected GM-RX, KIFD, LSUNRSORAD, and MD-RX as base methods for our supervised mGE-AD. We used these four base methods and thirty random channels from the HSI as input to the random forest meta-classifier. mGE-AD achieved an average of 0.841 ROC-AUC scores with a standard deviation (STD) of 0.213. Our baseline HUE-AD did not need any training. It is an averaging ensemble of AED, KIFD, LSUNRSORAD achieved an average of 0.854 ROC-AUC scores with a standard deviation (STD) of 0.251.
The individual results from the ABU dataset are shown as the airport in Table 4, the beach in Table 5, and the urban in Table 6. The Arizona dataset results are shown in Table 7.
The results from San Diego airport, HYDICE urban, and Salinas datasets are in Table 3. UGE-AD has the highest average in the HYDICE urban dataset. However, LSUNRSORAD and our baseline AD achieved a perfect score in the Salinas dataset. The baseline AD scores are better than others for the first scene of the San Diego airport dataset, whereas AED performed best in the second scene. The individual results from the ABU airport dataset are shown in Table 4, where AED performed best in the second and fourth scenes, LSUNRSORAD in the first scene, and baseline HUE-AD in the third scene. The individual results from the ABU beach dataset are shown in Table 5, where AED performed best in the first scene, baseline HUE-AD in the second scene, and UGE-AD in the fourth scene. However, most of the algorithms got perfect scores in the third scene. The individual results from the ABU urban dataset are shown in Table 6, where AED performed best in the first scene, baseline HUE-AD in the second scene, and UGE-AD in the fourth scene. However, most of the algorithms got perfect scores in the third scene.
Table 2: Average and standard deviation of ROC-AUC scores of various methods computed over the ABU, Arizona, HYDICE urban, Salinas, and San Diego datasets.
Methods Average Standard Deviation (STD)
CBAD 0.612 0.184
CSD 0.822 0.221
FCBAD 0.834 0.198
GM-RX 0.637 0.214
MD-RX 0.823 0.219
RX 0.826 0.222
SSRX 0.778 0.142
WIN-RX 0.769 0.154
AED 0.833 0.248
KIFD 0.887 0.141
LSUNRSORAD 0.833 0.269
baseline (AED, KIFD, LSUNRSORAD) 0.854 0.251
mGE-AD (GM-RX, KIFD, LSUNRSORAD, MD-RX) 0.841 0.213
UGE-AD (AED, FCBAD, GM-RX, KIFD) 0.918 0.131
Table 3: Comparison of ROC-AUC scores between various methods using the HYDICE urban, Salinas, and San Diego datasets
Methods HYDICE urban Salinas San Diego-01 San Diego-02
CBAD 0.970 0.390 0.674 0.547
CSD 0.982 0.788 0.941 0.871
FCBAD 0.994 0.841 0.969 0.958
GM-RX 0.957 0.546 0.557 0.219
MD-RX 0.986 0.817 0.945 0.900
RX 0.985 0.807 0.940 0.889
SSRX 0.747 0.848 0.768 0.836
WIN-RX 0.749 0.798 0.656 0.753
AED 0.979 0.956 0.985 0.991
KIFD 0.830 0.995 0.992 0.989
LSUNRSORAD 0.994 1.000 0.980 0.910
baseline HUE-AD 0.962 1.000 0.993 0.982
mGE-AD 0.927 0.997 0.916 0.962
UGE-AD 0.995 0.999 0.976 0.964
Table 4: Comparison of ROC-AUC scores between various methods using the ABU airport dataset
Methods abu-airport-1 abu-airport-2 abu-airport-3 abu-airport-4
CBAD 0.858 0.633 0.633 0.886
CSD 0.824 0.826 0.922 0.934
FCBAD 0.877 0.880 0.897 0.986
GM-RX 0.674 0.590 0.680 0.453
MD-RX 0.813 0.841 0.925 0.896
RX 0.822 0.840 0.929 0.953
SSRX 0.639 0.857 0.892 0.644
WIN-RX 0.774 0.797 0.949 0.708
AED 0.969 0.994 0.957 0.995
KIFD 0.940 0.979 0.961 0.975
LSUNRSORAD 0.972 0.959 0.966 0.959
baseline HUE-AD 0.970 0.989 0.973 0.988
mGE-AD 0.859 0.916 0.920 0.931
UGE-AD 0.933 0.971 0.948 0.986
Table 5: Comparison of ROC-AUC scores between various methods using the ABU beach dataset
Methods abu-beach-1 abu-beach-2 abu-beach-3 abu-beach-4
CBAD 0.571 0.603 0.664 0.885
CSD 0.982 0.911 1.000 0.956
FCBAD 0.987 0.955 1.000 0.948
GM-RX 0.986 0.701 1.000 0.916
MD-RX 0.980 0.910 1.000 0.955
RX 0.981 0.911 1.000 0.954
SSRX 0.980 0.906 0.982 0.831
WIN-RX 0.957 0.962 1.000 0.815
AED 0.997 0.955 1.000 0.979
KIFD 0.990 0.989 0.998 0.813
LSUNRSORAD 0.994 0.986 1.000 0.975
baseline HUE-AD 0.992 0.992 1.000 0.958
mGE-AD 0.962 0.980 0.981 0.935
UGE-AD 0.995 0.946 1.000 0.980
Table 6: Comparison of ROC-AUC scores between various methods using the ABU urban dataset
Methods abu-urban-1 abu-urban-2 abu-urban-3 abu-urban-4 abu-urban-5
CBAD 0.391 0.155 0.613 0.598 0.573
CSD 0.991 0.994 0.950 0.989 0.969
FCBAD 0.992 0.851 0.941 0.847 0.891
GM-RX 0.760 0.549 0.778 0.765 0.402
MD-RX 0.991 0.994 0.939 0.989 0.970
RX 0.991 0.995 0.951 0.989 0.969
SSRX 0.928 0.989 0.748 0.786 0.791
WIN-RX 0.927 0.909 0.783 0.787 0.881
AED 0.998 0.796 0.998 0.840 0.984
KIFD 0.938 0.917 0.992 0.978 0.984
LSUNRSORAD 0.996 0.972 0.988 0.928 0.983
baseline HUE-AD 0.993 0.972 0.996 0.985 0.990
mGE-AD 0.980 0.935 0.985 0.968 0.981
UGE-AD 0.977 0.976 0.962 0.973 0.952
Table 7: Comparison of ROC-AUC scores between various methods using the Arizona dataset
Methods Arizona 1 Arizona 2 Arizona 3 Arizona 4 Arizona 5
CBAD 0.510 0.374 0.556 0.742 0.648
CSD 0.335 0.410 0.601 0.264 0.645
FCBAD 0.367 0.432 0.662 0.350 0.727
GM-RX 0.576 0.271 0.457 0.760 0.423
MD-RX 0.334 0.410 0.602 0.271 0.648
RX 0.334 0.410 0.602 0.262 0.647
SSRX 0.564 0.578 0.661 0.458 0.682
WIN-RX 0.534 0.507 0.629 0.437 0.596
AED 0.329 0.326 0.492 0.307 0.504
KIFD 0.580 0.567 0.671 0.663 0.780
LSUNRSORAD 0.208 0.369 0.501 0.187 0.499
baseline HUE-AD 0.248 0.465 0.551 0.216 0.573
mGE-AD 0.427 0.290 0.622 0.382 0.639
UGE-AD 0.671 0.695 0.925 0.458 0.905
5 Conclusion
Our research showed that using different approaches does not always give similar results, as shown by the higher standard deviation (STD). Our new unsupervised stack ensemble method gives more reliable results with less variation, which is a positive development, and we plan to explore it further in future research.
Acknowledgements.
This research was sponsored by the Army Research Laboratory and was accomplished under Cooperative Agreement Number W911NF-21-2-0294 with Eddie L. Jacobs as PI. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the US Government. The US Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein. This material is based upon work supported by the National Science Foundation while Lan Wang was serving at the National Science Foundation. Any opinion, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.